OpenStudy (anonymous):

please help!!! Juan always saves the same amount from his weekly allowance. The table shows how much he has saved at different times. Which equation represents the situation? Juan’s Savings Weeks 3 6 9 12 Amount saved 19 28 37 46 A.y – 3 = −3(x – 19) B.y – 28 = 3(x – 6) C.y + 9 = −3(x + 37) D.y + 46 = 3(x – 12) i believe its d

OpenStudy (anonymous):

@abdullah1995 @amistre64

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Oops..!! you just missed the shot.. :(

OpenStudy (anonymous):

Equation of line with one point and slope is given as: \[y -y_1 = m(x-x_1)\]

OpenStudy (anonymous):

d would be right if it would have been : \(y - 46 = 3(x-12)\).. :)

OpenStudy (anonymous):

Any more try you wanna do?

OpenStudy (anonymous):

Slope is \(3\), that is right, so select equation with slope = 3 first.. :)

OpenStudy (anonymous):

so its b?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok hang on and thank you let me get another

OpenStudy (anonymous):

Point is \((6, 28)\) and \(m=3\), so: \(y - 28 = 3(x-6)\)

OpenStudy (anonymous):

Post it here only.. :

OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. A. (−6, −14) B. (−1, −7) C. (4, 10) D. (7, 9) i believe a,c,and d

OpenStudy (anonymous):

I am not seeing it, I am seeing boxes there.. :) Post it again and nicely.. :)

OpenStudy (anonymous):

What is the sign between 7x and 5y ?? Is it plus or minus?? Secondly, write A and B options again,,,

OpenStudy (anonymous):

\(a\) is right.. Why you chose \(c\) and \(d\) ?

OpenStudy (anonymous):

Just put that values in place of x and y, and check whether you are getting 28 or not.. :)

OpenStudy (anonymous):

Like for \(a\), (-6,-14) \[7(-6) -5(-14) = -42 + 70 = 28\] yes you are getting \(28\) here, so this is right.. Likewise do for others and check.. :)